OpenStudy (anonymous):

Can you answer this? Express the square of 3d^2+2d-5 as a polynomial

OpenStudy (anonymous):

No I can't. But thank you for asking.

OpenStudy (zehanz):

Square it like (a+b)²= a²+2ab+b²:\[(3d^2+2d-5)^2=(3d^2+(2d-5))^2\]Now a=3d² and b=2d-5, so you get:\[(3d^2)^2+2 \cdot3d^2 \cdot (2d-5) +(2d-5)^2\]Try to work out each of these three factors: \[(3d^2)^2=\]\[2 \cdot 3d^2(2d-5)=\]\[(2d-5)^2=\]For the last one, you need (a-b)^2=a²-2ab+b². It is a lot of work, but it is rewarding ;)